OpenStudy (anonymous):

SMART PEOPLE WHO ARENT BUSY AND WILLING TO HELP A HOME GIRL OUT.. REPLY TO ME! IM IN DIRE NEED!!!

OpenStudy (anonymous):

THEN QUIT SHOUTING AND ASK THE QUESTION :)

OpenStudy (anonymous):

i have been lol

OpenStudy (anonymous):

but anyways find a polynomial function whose graph passes thru (5,11)(10,-12)(0,3)

OpenStudy (anonymous):

We can find a quadratic polynomial of the form y = ax^2 + bx + c. First, since (x=0, y=3) is on the curve, if we plug these into the equation we get 3 = a(0^2)+b(0) +c, and thus c = 3. So we have so far y = ax^2 + bx + 3. If you write out the other equations with the remaining coordinates, we now have 2 equations with two unknows: 11 = 25a+5b+3 and -12 = 100a+10b+3. There are different methods for solving this, but once you do, you'll have a, b and c, and thus the equation of the quadratic.

OpenStudy (anonymous):

hi thanx